### v2

@@ 3 The Data Sample
Due to the diverse nature of various GRB classes, it is essential to avoid mixing the characteristics of these classes in an ML analysis. In a previous paper, some of us (Dainotti et al., 2024c) have pointed out that the removal of some additional classes, especially the SEE, does not significantly affect the prediction. Thus, this study focuses only on LGRBs, the most numerous and homogeneous subset. Thus, all SGRBs are removed from our sample. Our data comprises of LGRBs observed in $\gamma$-rays and X-rays by the BAT and XRT telescopes onboard the Swift. They are found in the Third Swift-BAT GRB Catalogue (Lien et al., 2016), and for data from 2016 until 12th December 2023, we use the NASA Swift GRB Search Tool. This sample is an extension of the one presented in Dainotti et al. (2024c), which itself was taken from Dainotti et al. (2020) and Srinivasaragavan et al. (2020). For training our models, we have removed the GRBs that have photometric redshifts; an example is the famous high-$z$ GRB090429B. Each GRB has ten properties (features). Four features belong to the prompt: the energy fluence over $T_{90}$ of the prompt emission in units of erg cm-2 (Fluence), $T_{90}$, the prompt peak photon flux in units of the number of photons cm-2 s-1 (Peak Flux), and from the BAT Telescope, the prompt photon index of the photon energy distribution modeled with a powerlaw (PhotonIndex). Six features belong to the X-ray plateau and afterglow: the time at the end of the plateau emission ($T_{a}$), its corresponding flux ($F_{a}$), the temporal powerlaw index after the end of the plateau emission ($\alpha$), the spectral index of the plateau ($\beta$) assuming a powerlaw for the spectral energy distribution, the neutral hydrogen column density along the line of sight (NH), the spectral index obtained as the time-averaged spectral fit from the Swift XRT Photon counting mode data ($\gamma$). The plateau emission feature is modeled using the so-called Willingale function (W07, Willingale et al. (2007)).
@@
Some of the parameters such as the $Peakflux$, $T_{a}$, $F_{a}$, and $T_{90}$, are shown the Fig. 1, with $T_{a}$, $F_{a}$ representing the x and y coordinates of the black dot. All parameters of the afterglow were obtained in 0.3-10 keV.
@@
The initial dataset contains 251 GRBs, including all classes. Compared with Dainotti et al. (2024c), there are an additional 30 GRBs with redshift that are added to the sample, an increase of 13%. From this sample, we remove 13 SGRBs and are left with 238 LGRBs with all the features listed above, as well as our response variable, the $\log(z+1)$ [^1] Note that all the logarithms in this work are in base 10. The response variable is the variable that we train our ML models to predict. We call this set of 238 GRBs the training set. This training set is 20% larger than the previous training set of Dainotti et al. (2024c). For the generalization set, we use the 221 GRBs of Dainotti et al. (2024c) and add 78 GRBs (an increase of 35%) taken from the Swift website. The additional 78 GRBs of the generalization set were fitted with the W07 model to obtain their $T_{a}$, $F_{a}$, $\alpha$, and $\beta$.
@@
Figure 1: This figure depicts a typical GRB LC. The X-ray LCs are fitted with the W07 model, denoted with the darker red line. The end time of the plateau ($T_{a}$,$F_{a}$) and the prompt emission peak are shown as a black-filled dot and as an empty dot, respectively.
@@
In Fig.2 we present the scatter matrix plot between the training set from Dainotti et al. (2024c) (cyan) and the 30 new GRBs (orange) added in this analysis for comparing the distribution of the variables. We applied the Kolmogorov-Smirov test (KS-test) and the Anderson-Darling test for each variable between the two data sets.
@@
All the variables pass the two tests with a p-value $>$ 0.05, indicating that they originate from the same parent population, except $T_{90}$ and $\gamma$, which obtain p-values $<$ 0.05 for both KS-test and AD-test, and $\log(F_{a})$ which obtained a p-value $<$ 0.05 for AD-test. For further discussion see the Appendix and the Fig. 19. Here, we are presenting the comparison between the redshift distributions.
@@
In the left panel of Fig. 3, the redshift distribution is illustrated, with the new sample depicted by a dashed line and the old sample by a solid line, both normalized to 1. The right panel of Fig. 3 presents the actual distribution with no normalization. The redshift distributions pass the KS-test with a p-value of 0.35 and the AD-test with p-value of 0.41, as it can be seen from the right panel of Fig. 3 . Thus, we cannot reject the hypothesis that the two sets of redshifts originate from the same parent population.
@@
Figure 2: The scatter matrix plot shows the distribution of the features involved, with the diagonal showing the distributions. Here, the additional 30 GRBs added to the training set are orange dots, while the previous sample from Dainotti et al. (2024c) is represented by cyan dots. The numbers in the upper triangle show the correlation between two respective features, measured for new GRBs, the previous sample, and the combined sample.
@@
Figure 3: Left panel: The redshift distribution of the new sample plotted in dashed and the old sample in continuous line normalized to 1. Right panel: the distribution of the new and the old samples not normalized.
@@ 4 Methodology and web app
In this research, we expand the methodology implemented in Dainotti et al. (2024c), and in addition, we make it available to the community via an easy-to-use and open-source web app. This web app is designed to let users input GRB parameters and quickly obtain pseudo-redshifts within hours of the GRB trigger. This will allow the community to conduct follow-up observations of interesting bursts, such as high-$z$ events or low-$z$ low-luminosity GRB (LLGRB) events. This web app is also highly modular, and users are free to train their own custom models using their data.
We here detail in each subsection the various methods implemented in our analysis and the associated functions of the web-app. The webapp is available for free on the following Github page:https://github.com/gammarayapp/GRB-Web-App
@@ 4.1 MICE
Figure 4: The figure shows the distribution of missing data in our data set. The figure is divided into rows and columns. The top axis of the columns corresponds to the features in the data set, and the bottom axis corresponds to the total number of missing entries in that specific feature. Thus, for $\log(NH)$ there are 19 missing entries, for $\log(Peakflux)$ there are 5 missing entries, and so on. The left axis corresponds to the number of GRBs, and the right axis corresponds to the number of features with missing entries. The blue squares show the complete entries, and the red squares show which features are missing for a given number of GRBs. Thus, the first row indicates that there are 209 GRBs with no missing features, and the second row indicates that there are 15 GRBs with missing entries in only 1 feature, which is $\log(NH)$. The third row indicates that there are 3 GRBs with missing entries in 1 feature, which is $\log(Peakflux)$, and the fourth row indicates that there are 2 GRBs with missing entries in 2 features simultaneously, which are $\log(Peakflux)$ and $\log(NH)$.
@@
Similar to the sample in Dainotti et al. (2024c), both our training and generalization sets have missing data. Out of the 238 LGRBs composing our training set, 29 GRBs have missing variables: 15 GRBs have missing $\log(NH)$ entries, 3 GRBs have missing $\log(Peakflux)$ entries, 2 GRBs have missing entries in both $\log(Peakflux)$ and $\log(NH)$ simultaneously, 3 GRBs have missing entries in $\alpha$, 2 GRBs have missing entries in $\alpha$ and $\log(NH)$ simultaneously, 2 GRBs have missing entries in $PhotonIndex$, 1 GRB has a missing entry in $\beta$ only, and 1 GRB has missing entries both $\gamma$ and $\beta$, simultaneously (see Fig. 4).
@@
In Dainotti et al. (2024c), we removed the GRBs from the generalization set with missing entries. However, in this analysis, we impute them in a similar fashion to the training set. It should be noted that the imputation for the training set and the generalization set are conducted separately so that the training set does not influence the generalization set, thus preventing data bleeding. All 299 GRBs in our generalization set are LGRBs, out of which 11 have missing data entries. Two GRBs have missing entries in $\gamma$, two have missing entries in $\log(NH)$, two GRBs have missing entries in both $\gamma$ and $\log(NH)$ simultaneously, two GRBs have missing entries in $\log(Peakflux)$ and $\log(T_{90})$ and three GRBs have simultaneous missing data in $PhotonIndex$, $\log(T_{90})$ and $\log(Peakflux)$. In addition to the previously noted missing entries, we identify GRBs with $\log(NH)<20$, $\alpha$, $\beta$, and $\gamma>3$, and a $PhotonIndex<0$ as outliers, as these values fall towards the extreme ends of their respective feature distributions. Instead of removing these values from the analysis, we consider these as missing data and impute them alongside the previously mentioned missing data.
@@
To minimize data loss and enhance the size of both the training and generalization sets, we utilized the imputation technique known as Multivariate Imputation by Chained Equations (MICE, Van Buuren & Groothuis-Oudshoorn (2011); Gibson et al. (2022)). MICE leverages the data currently in use to infer the missing data. Here, for the data imputation to be reliable, MICE implicitly assumes that the data is missing at random. This means that the observed data can explain the missing data, and the probability of the data not being observed does not depend on the missing data’s value itself. Furthermore, MICE also assumes that the data have a multivariate distribution and, thus, considers the relations between features for imputing missing data. Here, we stress that random imputation is appropriate, as the missing values of the variable are mainly due to the good and bad time intervals, orbital gaps, instrumental errors, and satellite follow-up repointing.
@@
The imputed data are sampled from the observed data, and the distribution is kept the same. We did this on purpose so that we would not have differences that could bias the analysis. In this study, we used the predictive mean matching method (PMM), specifically the mida-touch approach (see Little & Rubin (2019) for details). Using PMM ensures that the imputed data is sampled from the observed data. This method fills in a feature’s missing values with its mean and then iteratively refines these values by training a model on the complete data available. Each prediction is assigned a probability based on its proximity to the imputed value for the missing variable. The missing entry is then set by randomly choosing it from the observed values of the respective variable, weighted according to the previously defined probability. This process is carried out iteratively for each feature with missing entries until the values stabilize or a predetermined number of iterations is reached. We perform this iteration 20 times to have a reliable imputation. With the application of MICE, we increase our training sample by 14% and the generalization set by 4%.
@@
We also implement MICE imputation in the webapp, so that any user can upload data with missing entries and obtain a complete data set with imputed values. The web app also shows the missing entries in the user’s data, as shown in Fig. 4 when the ’Apply MICE’ option is selected (see Fig. 8).
@@ 4.2 Machine Learning methods
In the upcoming sections, we outline this study’s machine learning methods and models.
@@ 4.2.1 LASSO feature selection
Figure 5: LASSO coefficients for all 10 predictors. The top 7 are chosen for further analysis.
@@
The Least Absolute Shrinkage and Selection Operator (LASSO) is a method for selecting the best features for a given response variable. This method employs shrinkage for linear regression by constraining the $\ell^{1}$ norm (the sum of the magnitudes of all vectors in the given space) of the solution vector (in this case, $\log(z+1)$) to be less than or equal to a positive value called the tuning parameter ($\lambda$). This penalization enables the model to choose a subset of features and eliminate the remaining by assigning a zero coefficient (Tibshirani, 1996). In this work, we apply the LASSO selection feature using the GLMNET function (Hastie et al., 2017; Tibshirani et al., 2012; Dainotti et al., 2021a). Then, we select the features’ coefficients corresponding to the largest value of $\lambda$, with errors less than 1 standard deviation (Friedman et al., 2010). The top seven features, $\log(NH),\log(Peakflux),PhotonIndex,\log(T_{a}),\log(F_{a}),\gamma$, and $\alpha$, (see Fig.5) are picked for further analysis.
4.2.2 The ML Models used
@@
ML models in general require a large amount of data to train. Here, even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the amount of data required to reliably train more complex ML models, such as neural networks. Thus, we tested multiple different ML models to find the best set of models that can give accurate redshift predictions with the data set currently available. The three chosen models performed the best during our analysis. Furthermore, to test whether the performance of these models is maintained as the dataset increases, we have performed a Monte Carlo simulation as well (see Sec.4.7). Thus, based on our analysis we are confident that these chosen models are best suited for the data at hand. .
@@
  • Generalized Additive Model (GAM, Hastie & Tibshirani (1990)) is a semi-parametric method in which the response variable is connected to the predictors via the sum of either a parametric or non-parametric function or a combination of both. GAM provides the option for fitting smooth spline functions to the data that can enable non-linear fits. We explore the performance of GAM with and without the application of smoothing functions.
@@
  • Generalized Linear Model (GLM, Nelder & Wedderburn (1972)) is an extension of the linear model that uses an iteratively weighted least squares method for the estimation of maximum likelihood to fit the relation between the response variable and the predictors. GLM allows for non-linear distributions of the response variable through a link function. In our analysis, we use the default Gaussian link function.
@@
  • Random forest (RF) is tree based ML method that generates multiple independent regression trees (Breiman, 2001; Miller et al., 2015; Dainotti et al., 2021a). The final results are an average of these regression trees. RF has three hyperparameters: depth, number of trials, and number of trees. Hyperparameters are function parameters that determine a function’s properties and dictate how well they fit a given set of data. In our analysis, we use the default implementation of RF from the Caret package (Kuhn & Max, 2008).

### v3

@@ 3 The data sample
Due to the diverse nature of various GRB classes, it is essential to avoid mixing the characteristics of these classes in an ML analysis. In a previous paper (Dainotti et al., 2024c), some members of our research team have pointed out that the removal of some additional classes, especially the SEE, does not significantly affect the prediction. Thus, this study focuses only on LGRBs, the most numerous and homogeneous subset, and all SGRBs were removed from our sample. Our data comprises of LGRBs observed in $\gamma$-rays and X-rays by the BAT and XRT telescopes onboard the Swift. They are found in the Third Swift-BAT GRB Catalogue (Lien et al., 2016), and for data from 2016 until 12th December 2023, we use the NASA Swift GRB Search Tool. This sample is an extension of the one presented in Dainotti et al. (2024c), which itself was taken from Dainotti et al. (2020) and Srinivasaragavan et al. (2020). For training our models, we have removed the GRBs that have photometric redshifts; an example is the famous high-$z$ GRB090429B. Each GRB has ten properties (features). Four features belong to the prompt: the energy fluence over $T_{90}$ of the prompt emission in units of erg cm-2 (Fluence), $T_{90}$, the prompt peak photon flux in units of the number of photons cm-2 s-1 (Peak Flux), and from the BAT Telescope, the prompt photon index of the photon energy distribution modeled with a power law (PhotonIndex). Six features belong to the X-ray plateau and afterglow: the time at the end of the plateau emission ($T_{a}$), its corresponding flux ($F_{a}$), the temporal power-law index after the end of the plateau emission ($\alpha$), the spectral index of the plateau ($\beta$) assuming a power law for the spectral energy distribution, the neutral hydrogen column density along the line of sight (NH), the spectral index obtained as the time-averaged spectral fit from the Swift XRT Photon counting mode data ($\gamma$). The plateau emission feature is modeled using the so-called Willingale function (W07, Willingale et al. (2007)).
@@
Some of the parameters such as the $Peakflux$, $T_{a}$, $F_{a}$, and $T_{90}$, are shown the Fig. 1, with $T_{a}$, $F_{a}$ representing the x and y coordinates of the black dot. All of the parameters of the afterglow were obtained in 0.3-10 keV.
@@
The initial data set contains 251 GRBs, including all classes. Compared with Dainotti et al. (2024c), there are an additional 30 GRBs with redshift that are added to the sample, an increase of 13%. From this sample, we remove 13 SGRBs and are left with 238 LGRBs with all the features listed above, as well as our response variable, the $\log(z+1)$ [^1] We note that all the logarithms in this work are in base 10. The response variable is the variable that we train our ML models to predict. We call this set of 238 GRBs the training set. This training set is 20% larger than the previous training set of Dainotti et al. (2024c). For the generalization set, we use the 221 GRBs of Dainotti et al. (2024c) and add 78 GRBs (an increase of 35%) taken from the Swift website. The additional 78 GRBs of the generalization set were fit with the W07 model to obtain their $T_{a}$, $F_{a}$, $\alpha$, and $\beta$.
@@
Figure 1: Typical GRB LC. The X-ray LCs are fit with the W07 model, denoted with the darker red line. The end time of the plateau ($T_{a}$,$F_{a}$) and the prompt emission peak are shown as a black-filled dot and as an empty dot, respectively.
@@
In Fig.2 we present the scatter matrix plot between the training set from Dainotti et al. (2024c) (cyan) and the 30 new GRBs (orange) added in this analysis for comparing the distribution of the variables. We applied the Kolmogorov-Smirov test (KS test) and the Anderson-Darling test (AD test) for each variable between the two data sets.
@@
All the variables pass the two tests with a p-value $>$ 0.05, indicating that they originate from the same parent population, except $T_{90}$ and $\gamma$, which obtain p-values $<$ 0.05 for both KS test and AD test, and $\log(F_{a})$ which obtained a p-value $<$ 0.05 for AD test. For further discussion see the Appendix and the Fig. 19. Here, we are presenting the comparison between the redshift distributions.
@@
In the left panel of Fig. 3, the redshift distribution is illustrated, with the new sample depicted by a dashed line and the old sample by a solid line, both normalized to one. The right panel of Fig. 3 presents the actual distribution with no normalization. The redshift distributions pass the KS test with a p-value of 0.35 and the AD test with p-value of 0.41, as it can be seen from the right panel of Fig. 3 . Thus, we cannot reject the hypothesis that the two sets of redshifts originate from the same parent population.
@@
Figure 2: Scatter matrix plot showing the distribution of the features involved, with the diagonal showing the distributions. The 30 GRBs added to the training set are shown as orange dots, while those of the previous sample from Dainotti et al. (2024c) are represented by cyan dots. The numbers in the upper triangle show the correlation between two respective features measured for new GRBs, the previous sample, and the combined sample.
@@
Figure 3: Left panel: Redshift distribution of the new sample (dashed lines) and the old sample (solid line) normalized to one. Right panel: Distribution of the new and the old samples when not normalized.
@@ 4 Methodology and web app
In this research, we expand the methodology implemented in Dainotti et al. (2024c), and in addition, we make the webapp available to the community via an easy-to-use open-source web app. This web app is designed to let users input GRB parameters and quickly obtain pseudo-redshifts within hours of the GRB trigger. This will allow the community to conduct follow-up observations of interesting bursts, such as high-$z$ events or low-$z$ low-luminosity GRB (LLGRB) events. This web app is also highly modular, and users are free to train their own custom models using their data.
We here detail in each subsection the various methods implemented in our analysis and the associated functions of the web app.
@@ 4.1 MICE
Figure 4: Distribution of missing data in our data set. The figure is divided into rows and columns. The top axis of the columns corresponds to the features in the data set, and the bottom axis corresponds to the total number of missing entries in that specific feature. Thus, for $\log(NH)$ there are 19 missing entries, for $\log(Peakflux)$ there are five missing entries, and so on. The left axis corresponds to the number of GRBs, and the right axis corresponds to the number of features with missing entries. The blue squares show the complete entries, and the red squares show which features are missing for a given number of GRBs. Thus, the first row indicates that there are 209 GRBs with no missing features, and the second row indicates that there are 15 GRBs with missing entries in only one feature, which is $\log(NH)$. The third row indicates that there are three GRBs with missing entries in one feature, which is $\log(Peakflux)$, and the fourth row indicates that there are two GRBs with missing entries in two features simultaneously, which are $\log(Peakflux)$ and $\log(NH)$.
@@
Similar to the sample in Dainotti et al. (2024c), both our training and generalization sets have missing data. Out of the 238 LGRBs composing our training set, 29 GRBs have missing variables: 15 GRBs have missing $\log(NH)$ entries, three GRBs have missing $\log(Peakflux)$ entries, two GRBs have missing entries in both $\log(Peakflux)$ and $\log(NH)$ simultaneously, three GRBs have missing entries in $\alpha$, two GRBs have missing entries in $\alpha$ and $\log(NH)$ simultaneously, two GRBs have missing entries in $PhotonIndex$, one GRB has a missing entry in $\beta$ only, and one GRB has missing entries in both $\gamma$ and $\beta$ simultaneously. (see Fig. 4).
@@
In Dainotti et al. (2024c), we removed the GRBs from the generalization set with missing entries. However, in this analysis, we impute them in a similar fashion to the training set. It should be noted that the imputation for the training set and the generalization set are conducted separately so that the training set does not influence the generalization set, thus preventing data bleeding. All 299 GRBs in our generalization set are LGRBs, out of which 11 have missing data entries. Two GRBs have missing entries in $\gamma$, two have missing entries in $\log(NH)$, two GRBs have missing entries in both $\gamma$ and $\log(NH)$ simultaneously, two GRBs have missing entries in $\log(Peakflux)$ and $\log(T_{90})$ and three GRBs have simultaneous missing data in $PhotonIndex$, $\log(T_{90})$ and $\log(Peakflux)$. In addition to the previously noted missing entries, we identify GRBs with $\log(NH)<20$, $\alpha$, $\beta$, and $\gamma>3$, and a $PhotonIndex<0$ as outliers, as these values fall toward the extreme ends of their respective feature distributions. Instead of removing these values from the analysis, we consider these as missing data and impute them alongside the previously mentioned missing data.
@@
To minimize data loss and enhance the size of both the training and generalization sets, we utilized the imputation technique known as multivariate imputation by chained equations (MICE; Van Buuren & Groothuis-Oudshoorn (2011); Gibson et al. (2022)). MICE leverages the data currently in use to infer the missing data. Here, for the data imputation to be reliable, MICE implicitly assumes that the data are missing at random. This means that the observed data can explain the missing data, and the probability of the data not being observed does not depend on the missing data’s value itself. Furthermore, MICE also assumes that the data have a multivariate distribution and thus considers the relations between features for imputing missing data. Here, we stress that random imputation is appropriate, as the missing values of the variable are mainly due to the good and bad time intervals, orbital gaps, instrumental errors, and satellite follow-up repointing.
@@
The imputed data are sampled from the observed data, and the distribution is kept the same. We did this on purpose so that we would not have differences that could bias the analysis. In this study, we used the predictive mean matching method (PMM), specifically the mida-touch approach (see Little & Rubin (2019) for details). Using PMM ensures that the imputed data are sampled from the observed data. This method fills in a feature’s missing values with its mean and then iteratively refines these values by training a model on the complete data available. Each prediction is assigned a probability based on its proximity to the imputed value for the missing variable. The missing entry is then set by randomly choosing it from the observed values of the respective variable, weighted according to the previously defined probability. This process is carried out iteratively for each feature with missing entries until the values stabilize or a predetermined number of iterations is reached. We perform this iteration 20 times to have a reliable imputation. With the application of MICE, we increase our training sample by 14% and the generalization set by 4%.
@@
We also implemented MICE imputation in the web app so that any user can upload data with missing entries and obtain a complete data set with imputed values. The web app also shows the missing entries in the user’s data, as shown in Fig. 4 when the ”Apply MICE” option is selected (see Fig. 8).
@@ 4.2 Machine learning methods
In the upcoming sections, we outline this study’s machine learning methods and models. We discuss the feature selection methodology, the ML models used and the formula generation methodology.
@@ 4.2.1 LASSO feature selection
Figure 5: LASSO coefficients for all ten predictors. The top seven were selected for further analysis.
@@
The least absolute shrinkage and selection operator (LASSO) is a method for selecting the best features for a given response variable. This method employs shrinkage for linear regression by constraining the $\ell^{1}$ norm (the sum of the magnitudes of all vectors in the given space) of the solution vector (in this case, $\log(z+1)$) to be less than or equal to a positive value called the tuning parameter ($\lambda$). This penalization enables the model to choose a subset of features and eliminate the remaining by assigning a zero coefficient (Tibshirani, 1996). In this work, we apply the LASSO selection feature using the GLMNET function (Hastie et al., 2017; Tibshirani et al., 2012; Dainotti et al., 2021a). Then, we select the features’ coefficients corresponding to the largest value of $\lambda$, with errors less than one standard deviation (Friedman et al., 2010). The top seven features, $\log(NH),\log(Peakflux),PhotonIndex,\log(T_{a}),\log(F_{a}),\gamma$, and $\alpha$, (see Fig.5) are picked for further analysis.
4.2.2 The machine learning models used
@@
The ML models in general require a large amount of data to train. Here, even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the amount of data required to reliably train more complex ML models, such as neural networks. Thus, we tested multiple different ML models to find the best set of models that can give accurate redshift predictions with the data set currently available. The three chosen models performed the best during our analysis. Furthermore, to test whether the performance of these models is maintained as the data set increases, we have performed a MC simulation as well (see Sec.4.7). Thus, based on our analysis we are confident that these chosen models are best suited for the data at hand.
@@
  • Generalized additive model (GAM; Hastie & Tibshirani (1990)) is a semi-parametric method in which the response variable is connected to the predictors via the sum of either a parametric or non-parametric function or a combination of both. GAM provides the option for fitting smooth spline functions to the data that can enable non-linear fits. We explore the performance of GAM with and without the application of smoothing functions.
@@
  • Generalized linear model (GLM; Nelder & Wedderburn (1972)) is an extension of the linear model that uses an iteratively weighted least squares method for the estimation of maximum likelihood to fit the relation between the response variable and the predictors. GLM allows for non-linear distributions of the response variable through a link function. In our analysis, we use the default Gaussian link function.
@@
  • Random forest (RF) is a tree-based ML method that generates multiple independent regression trees (Breiman, 2001; Miller et al., 2015; Dainotti et al., 2021a). The final results are an average of these regression trees. RF has three hyperparameters: depth, number of trials, and number of trees. Hyperparameters are function parameters that determine a function’s properties and dictate how well they fit a given set of data. In our analysis, we use the default implementation of RF from the Caret package (Kuhn & Max, 2008).